# CHIRLA: Comprehensive High-resolution Identification and Re-identification for Large-scale Analysis

### v1

@@ Methods
Train and Test Data
This Section provides a comprehensive overview of the training and test datasets defined for each tracking and Re-ID benchmark scenario. It details the data format, organizational structure, and the methodology employed for data generation and annotation.
Tracking Data
For the tracking benchmark, the train set consists of two sequences, while the test set includes eight sequences. This distribution remains consistent across the two tracking scenarios defined:
  • Tracking after brief occlusions. Brief occlusions are defined as interruptions lasting between one and five seconds. For this scenario, only the frames immediately before and after the occlusion starts and ends are included for each ID, along with their corresponding bounding box coordinates.
    The data is structured in a hierarchical JSON format. Each JSON file consists of a dictionary where the keys represent individual IDs in the video, and the values are lists of annotations. Each annotation is represented as a dictionary containing four keys:
    • "frame_end": The frame number immediately after the occlusion ends.
  • Tracking in scenarios with multiple people and occlusions. This scenario extends the conditions of the brief occlusions scenario by requiring that the occluded ID has an intersection with another ID before the occlusion begins. An IoU threshold of 0.3 is applied to ensure sufficient overlap between bounding boxes before the ID is occluded. The data for multiple occlusions is represented using the same JSON format as in the brief occlusions scenario.
Re-Identification Data
Person Re-ID experiments focus on evaluating frames following long occlusions, which are defined as lasting more than five seconds. This distinction ensures a clear differentiation from tracking scenarios that involve shorter occlusions.
Across the four scenarios defined in the Re-ID benchmark, the training sets include the first visible frame of each ID, followed by four consecutive frames sampled at an 18-frame gap (equivalent to 2.4 seconds of visibility). In contrast, the test sets comprise the reappearance frames and the four subsequent frames for each reappearance, using the same sampling strategy as the training sets. However, a minimum of 20 frames per ID is required in the test sets. If fewer than 20 frames are available, additional frames are randomly sampled to meet the requirement.
In addition, IDs present in the dataset are assigned negative values in the test sets when they meet the long-term occlusion conditions but are not included in the training sets. This labeling explicitly defines them as unknown IDs, providing an additional challenge to the Re-ID system.
All data is organized in a hierarchical JSON structure. Each JSON file consists of a dictionary where the keys represent individual IDs within the video, and the values are lists of annotations. Each annotation is represented as a dictionary with the following two keys:
  • "frame": The frame number where the person is present in the video.
For each benchmark scenario, subsets of training and test data are provided. The final evaluation of benchmark results is computed as the average of the results from individual subsets. The specific details of the training and test data for each Re-ID scenario are outlined below. In all cases, the videos selected for evaluation were carefully analyzed to ensure they align with the benchmark definitions for each Re-ID task, with an emphasis on presenting challenging cases. Figure 5 illustrates sample data for one of the Re-ID scenario benchmarks.
  • Re-identification after reappearance on the scene. This scenario includes 10 train and test set splits. The training set consists of annotations derived from 10 videos, with each subset containing one video per sequence. However, the selection was made from nine distinct sequences, as sequences that did not meet the long occlusion conditions were excluded. To complete the set, one sequence was repeated, sampling two different videos. If uninterrupted frames were not available within the analyzed duration (2.4 seconds), only the frames within the specified range were included. For each video analyzed in the training set, the corresponding test set includes all frames from the same video, where the individual reappears following a long occlusion.
  • Long-term re-Identification. The train and test sets consist of seven subsets. The training set includes annotations from one video per camera, resulting in a total of seven videos, with one video in each subset. The corresponding test subset comprises all remaining videos from the same camera analyzed in the training set, across different sequences.
  • Multi-Camera re-Identification. In this scenario, data is distributed across 10 training and test subsets. The training set consists of annotations from one video per sequence, resulting in 10 videos in total, with one video in each subset. The corresponding test subset includes all remaining videos within each analyzed sequence.
  • Multi-Camera Long-term re-Identification. In this case we also provide 10 subsets for training and testing. The training set consists of annotations from one video per sequence, resulting in 10 videos in total, with one video in each subset. Each test subset includes two videos from one or two sequences, which are always distinct from those used in the corresponding training subset.
    ((a)) Train data samples
    ((b)) Test data samples
Organization
The benchmark data is organized into two main directories: tracking and reid. Each directory contains subdirectories corresponding to the different scenarios, each further divided into training and test sets.
  • Tracking Directory Structure
    The tracking directory includes two subdirectories, one for each tracking benchmark scenario: brief_occlusions and multiple_people_occlusions. Each scenario is divided into train and test subdirectories, organized by sequence name (e.g., seq_004). Within each sequence subdirectory, the data is further structured by camera views. For each sequence, seven JSON files are provided, one per camera view, labeled with the camera name and metadata (e.g., camera_7_2023-06-08-12:29:08.json). Alongside the JSON files, there is an imgs directory, which contains seven subdirectories, one for each camera view. These subdirectories share the same names as the corresponding JSON files.
    Inside each camera view subdirectory, the data is organized into subfolders named after the IDs present in the corresponding video. Each ID subfolder contains cropped bounding box images for that ID, selected based on the defined tracking scenario. The images are stored in PNG format and are named as frame_n.png, where n corresponds to the frame number in the original video.
  • Re-identification Directory Structure
    The reid directory includes four scenarios: reappearance, long_term, multi_camera, and multi_camera _long_term. Each scenario has its own train and test subdirectories. Within each train and test directory, there are subset directories named as train_X or test_X, where X is the subset index (e.g., train_0, test_1). The long_term scenario contains seven subsets in both the training and test sets, while reappearance, multi_camera, and multi_camera_long_term contain 10 subsets for training and testing.
    Inside these directories, the structure follows the same format as the one outlined for the tracking benchmark, including sequences, JSON files, and imgs directories. The number of data points within each train and test set depends on the data splitting methodology described in Section Re-Identification Data.

### v2

@@ Methods
Benchmark Evaluation
In Table 3 we report the results of different tracking methods evaluated on the CHIRLA benchmark for person tracking under the two proposed scenarios. We first present the results on the full test set (8 sequences with a total of 56 videos) using the metrics described in Section Person Tracking Methods, which reflect the general performance of person tracking across the dataset. We then provide results for the two focused benchmark scenarios: tracking after brief occlusions and tracking with multiple people and occlusions.
Method All test set Brief occlusions Multiple people and occlusions
[…]
BoostTrack [32] 60.12 36.87 143 17 217654 89542 3801 7841 5.78 10.759 10 9.15 16.56 2
BoT-SORT [28] 56.52 35.33 245 6 102056 232112 4883 4991 12.42 19.15 91 21.831 33.88 10
ByteTrack [27] 56.10 34.96 241 6 106270 230720 5400 5705 11.57 17.77 87 20.42 31.87 11
OC-SORT [29] 62.86 34.55 208 9 137854 145336 6443 8434 10.71 17.10 51 15.50 25.89 6
BoostTrack+ [32] 60.12 37.50 146 17 216848 90364 3823 7884 6.63 11.91 11 9.86 17.72 2
BoostTrack++ [33] 62.14 39.62 145 21 230102 62099 3069 7951 5.78 10.71 13 11.27 20 2
BoT-SORT [28] 68.84 42.54 184 17 188089 52600 2298 4105 9.86 17.01 36 15.49 25.89 6
Deep-OC-SORT [30] 65.83 37.05 197 11 154621 106314 5575 8174 9.35 15.86 38 15.49 26.35 3
StrongSORT [31] 61.37 36.67 222 8 126838 166387 8054 8259 10.54 16.88 61 17.61 28.90 6
Table 3: Evaluation of different tracking methods on CHIRLA benchmark. Results on the full test set (8 sequences, 56 videos) are reported with MOTA, IDF1, MT/ML, FN, FP, IDSW and FM; the two focused scenarios (“Brief occlusions” and “Multiple people and occlusions”) report MOTA, IDF1 and IDSW. Arrows ($\uparrow$ / $\downarrow$) indicate whether higher or lower values are better. denotes models that incorporate appearance-embedding-based tracking. Appearance embeddings are extracted using a lightweight OSNet [40] backbone (×0.25 scaling) pretrained on MSMT17 dataset. For all methods, YOLOv11n [41] is used as the person detector, and all trackers are executed with the default parameters from their official repositories.
For both scenarios, after running the trackers on the full test set, we selected the ground-truth frames corresponding to each scenario (as defined in Section Tracking Benchmark Data Generation). For each ground-truth ID, we retained the predicted bounding box with the highest IoU (>0.5), along with the corresponding tracker ID. Using this mapping, we then computed the evaluation metrics to assess whether trackers consistently maintained ID assignment before and after the occlusion conditions.
On the full test set, BoT-SORT and its appearance-augmented variant BoT-SORT [28] achieved the best overall results. BoT-SORT reached the highest accuracy, with MOTA 68.84 and IDF1 42.54, while also yielding the lowest FP (52,600), IDSW (2,298) and FM (4105). Meanwhile, BoT-SORT achieved the highest number of MT trajectories (245) and the lowest FP (102,056), and together with ByteTrack [27] recorded the smallest ML count (6). This confirms that trackers integrating appearance analysis provide a clear advantage in general conditions with continuous observations.
However, in the two focused scenarios: brief occlusions and multiple people with occlusions; appearance-based trackers underperform compared to traditional trajectory-based methods. For example, in the brief occlusion scenario, BoT-SORT achieved the best performance (MOTA 12.42, IDF1 19.15), whereas BoostTrack variants [32, 33] reduced IDSW to 10 but struggled in MOTA. Similarly, in the multiple people and occlusions scenario, BoT-SORT again led (MOTA 21.83, IDF1 33.88), while BoostTrack variants minimized IDSW (2) but at the cost of MOTA and IDF1.
We believe this inversion of performance reflects the challenge of Re-ID after occlusions. Appearance embeddings are sensitive to even small changes caused by lighting, pose, or occlusion, making it difficult for appearance-based trackers to maintain consistent IDs. In contrast, traditional trackers, which rely mainly on motion and trajectory continuity, are sometimes better able to reassign IDs after short occlusions, without taking into account appearance cues.
Overall, results in both scenarios dropped significantly compared to the full test set, from MOTA 68.84 / IDF1 42.54 to 12.42 / 19.15 in brief occlusions and 21.83 / 33.88 in multiple people and occlusions. This performance gap illustrates the challenging nature of CHIRLA’s occlusion-driven benchmarks and highlights the need for more robust strategies that can combine motion continuity with Re-ID embeddings tolerant to appearance changes.
In addition, Table 4 reports the results of several models across the four CHIRLA Re-ID scenarios. Performance is evaluated using CMC at rank-1, rank-5 and rank-10, together with mAP, on the test sets defined for each scenario (see Section Train and Test Data). We restrict the experiments to closed-set evaluation, where only known IDs are included in both train and test sets. Specifically, distractors (unknown IDs) provided by CHIRLA are excluded from the test set. We leave the open-set evaluation for future studies, as it requires models to define decision thresholds to distinguish between known and unknown IDs, which would enable a more realistic assessment of deployment conditions.
For all scenarios, we exclude the subsets train_0 and test_0 from evaluation, as these can be reserved for potential fine-tuning or model adjustment. Thus, the training set is used as the gallery and the test set as the query. Since the dataset is organized into multiple train–test subsets, we run experiments on each train–test pair and report the final results as the average across all subsets.
Method Long term Multi-camera long term Multi-camera Reappearance
CMC mAP CMC mAP CMC mAP CMC mAP
@1 @5 @10 @1 @5 @10 @1 @5 @10 @1 @5 @10
ResNet50 [17] 13.30 33.19 49.39 18.22 18.02 36.92 61.58 30.65 42.58 60.02 75.00 45.39 50.98 75.30 89.92 54.98
ResNet50x[17] 11.64 33.47 49.57 16.96 16.55 37.51 59.18 28.50 36.21 57.02 69.97 40.02 52.04 76.91 89.35 54.36
ResNet50-IBN [18, 42] 4.54 22.04 40.36 12.84 11.12 37.30 63.93 23.30 13.24 45.21 64.21 25.82 19.61 50.39 74.90 31.36
ResNet50-IBNx [18, 42] 8.48 33.05 50.26 15.18 16.47 48.56 74.53 29.04 21.83 48.09 65.17 28.70 19.52 58.46 80.04 31.24
ResNet101-IBN [18, 43] 15.85 39.18 57.15 20.16 20.49 41.88 67.80 33.51 38.41 60.54 75.26 42.79 57.99 82.31 93.07 60.38
ResNet101-IBNx[18, 43] 11.75 34.02 52.74 17.43 16.32 43.65 64.88 29.18 35.14 59.98 74.53 39.62 56.08 80.14 91.87 57.36
ResNet101-IBN [18, 43] 12.22 32.81 51.14 18.45 22.48 52.63 73.93 35.28 53.45 72.79 83.45 53.25 61.00 83.29 93.00 61.15
ResNet101-IBN [18, 44] 12.61 36.48 56.19 19.64 19.88 43.33 68.68 32.08 39.43 59.15 80.30 43.97 58.37 82.12 94.38 60.59
ResNet101-IBNx [18, 44] 12.94 34.41 55.24 17.96 17.58 46.34 71.53 29.92 38.94 67.21 80.39 43.16 55.85 81.32 90.49 55.20
ResNet101-IBN [18, 44] 12.79 37.49 56.79 18.81 19.33 46.77 71.41 33.37 49.13 69.33 83.09 50.78 65.73 84.99 92.83 61.34
ResNet101-IBN [18, 45] 17.47 44.17 63.85 22.68 18.85 43.86 69.73 32.79 45.52 66.73 80.31 48.06 61.30 83.02 93.02 61.55
ResNet101-IBNx [18, 45] 13.36 38.42 58.43 18.51 18.52 44.36 70.68 28.72 38.45 61.11 75.43 40.98 55.59 80.34 90.73 54.96
ResNet101-IBN [18, 45] 17.68 43.44 64.06 21.46 20.02 49.48 74.18 30.53 54.54 73.24 85.58 52.61 61.29 83.93 92.91 59.39
ResNet50* [17, 42] 15.74 40.90 60.60 20.28 21.10 47.01 72.36 32.28 50.00 67.35 77.11 49.72 54.33 81.28 89.92 56.64
ResNet101* [17, 42] 18.58 43.76 62.07 22.01 24.31 45.79 67.18 33.93 50.70 68.33 77.50 49.97 56.88 82.01 92.20 58.69
ResNet152* [17, 42] 17.38 42.42 63.17 22.03 20.09 42.63 74.35 32.69 51.53 69.80 78.63 50.82 58.30 83.92 92.25 60.09
ResNet50-IBN* [18, 42] 8.61 26.10 43.39 15.79 14.40 38.81 58.15 27.78 23.61 47.67 68.64 33.78 43.64 65.90 84.57 46.33
ResNet101-IBN* [18, 42] 7.57 23.30 39.54 15.43 14.78 40.31 62.55 28.01 27.58 47.53 70.73 36.65 39.26 64.06 81.18 44.28
ResNet152-IBN* [18, 42] 7.68 25.22 39.58 15.41 18.14 40.23 62.89 30.24 22.96 45.00 68.00 34.16 41.47 67.58 84.28 46.94
ConvNext-B [19] 4.98 21.79 39.72 12.42 14.36 43.82 72.41 28.14 11.96 35.88 67.50 23.80 20.07 55.90 76.36 32.97
ConvNext-S [19] 11.48 38.90 59.62 16.26 21.85 47.68 69.75 30.77 14.55 39.57 67.11 25.01 17.97 50.53 71.47 30.06
ConvNext-T [19] 7.23 30.77 49.05 13.41 14.35 46.30 72.37 27.35 14.61 44.68 66.51 22.40 17.34 49.89 74.00 32.88
EdgeNext-B [20] 5.85 26.58 52.43 14.63 7.90 39.90 63.58 23.84 13.50 45.80 63.43 26.72 17.52 50.69 76.45 29.35
EdgeNext-S [20] 7.31 31.10 53.12 15.26 10.95 32.64 57.51 23.92 14.64 51.69 78.38 25.79 20.31 50.82 75.26 31.75
EdgeNext-XS [20] 9.11 28.54 39.94 14.60 17.33 50.35 70.46 29.89 12.24 47.90 64.40 24.34 17.17 46.17 77.39 30.59
FastViT-S12 [25] 7.99 27.85 53.15 15.45 13.00 42.04 71.73 26.57 12.85 41.00 59.95 23.86 26.17 51.69 74.02 33.81
FastViT-SA12 [25] 8.56 21.50 40.14 13.02 9.61 37.20 52.12 23.10 20.01 48.93 70.65 29.78 25.12 58.47 84.79 32.38
FastViT-SA24 [25] 8.48 20.60 38.78 12.50 7.26 45.41 75.18 24.07 19.48 44.83 67.22 28.44 19.31 47.05 74.23 30.48
[…]
RepViT-M0.9 [26] 3.88 36.87 55.48 13.62 8.39 52.29 74.03 23.96 20.30 53.57 74.96 26.21 17.35 61.88 81.19 28.12
RepViT-M1.0 [26] 7.22 31.38 60.07 13.55 10.26 60.97 85.51 23.46 16.36 56.48 86.48 25.30 22.85 70.39 88.45 29.55
RepViT-M1.5 [26] 7.46 33.93 48.08 13.40 16.69 46.54 87.65 25.60 10.72 53.37 86.49 21.91 18.32 68.29 83.71 27.49
ResNet50 [17] 14.23 39.28 59.72 20.20 17.74 43.32 72.81 31.27 52.16 70.64 79.71 53.31 55.06 81.71 91.49 58.57
ResNet101 [17] 18.81 44.27 63.25 23.24 19.32 47.82 69.99 33.24 53.46 70.02 78.37 54.26 62.55 85.53 93.04 62.88
ResNet152 [17] 18.21 43.98 62.84 23.04 21.11 44.33 75.55 33.85 56.01 72.35 81.10 55.90 63.59 86.22 93.03 63.85
ResNet18-IBN [18] 9.41 25.64 40.73 16.25 13.91 38.75 61.93 26.91 20.95 47.54 66.69 31.97 30.89 59.35 83.32 40.71
ResNet50-IBN [18] 8.09 27.09 43.70 15.67 15.38 39.90 57.85 27.86 21.99 46.54 68.36 32.61 43.09 65.79 83.19 45.86
ResNet101-IBN [18] 8.54 24.16 39.68 15.71 16.05 40.79 62.88 28.91 26.68 49.20 69.69 36.45 38.87 62.46 80.75 44.03
ResNet152-IBN [18] 7.74 24.99 39.15 15.37 17.59 42.32 64.84 30.32 23.60 45.28 67.16 34.12 40.46 68.05 84.42 46.78
Swin-S [23] 7.02 25.81 52.77 14.88 16.18 51.18 74.66 27.51 19.77 50.90 73.58 29.49 16.94 47.38 75.47 30.40
Swin-T [23] 11.22 43.75 61.95 15.20 11.74 46.62 76.67 25.71 14.04 41.01 67.33 27.15 16.88 51.21 77.45 30.98
ViT-S [22] 6.11 27.12 43.88 14.73 15.84 40.77 58.55 27.17 15.93 38.44 66.55 29.76 20.15 42.20 66.64 32.39
ViT-T [22] 7.14 29.19 45.85 15.05 16.31 38.08 67.90 26.59 14.63 42.32 63.73 29.94 19.44 40.69 66.40 32.38
VOLO-D1 [24] 12.66 34.09 53.90 17.95 19.50 42.41 70.24 31.80 38.57 58.23 71.18 40.95 45.59 73.17 88.45 49.98
VOLO-D2 [24] 12.35 34.50 50.98 17.59 17.46 44.24 69.91 31.10 40.02 59.66 70.94 41.78 49.70 75.03 85.68 49.97
VOLO-D3 [24] 13.99 36.84 54.40 18.94 20.36 48.81 70.59 30.41 35.33 64.93 77.21 40.99 53.83 81.97 92.54 54.11
Table 4: Evaluation of the four person re-identification scenarios defined in our proposed benchmark using different methods in closed-set evaluation. indicates specific pre-training on Market1501, x on DukeMTMC, and on MSMT17 datasets. * denotes results with an input size of $384\times 128$, while all other results use $256\times 128$. Model weights for are extracted from CentroidsReid [46], for from FastReid [45], and the remaining models from ReIDZoo [47].
The long-term scenario emerges as the most challenging, as it requires consistent Re-ID over extended sequences with large temporal gaps. This is reflected in the lowest CMC values among all scenarios. The best model here is ResNet101 [17] (256×128 input size), which reached 18.81% CMC@1 and 23.24% mAP, highlighting the difficulty of maintaining reliable embeddings across long time spans.
The multi-camera long-term scenario results as the second most difficult. In this case, a ResNet101 backbone with a larger input size of 384×128 achieved the best CMC@1 result (24.31%), while ResNet101-IBN [18] obtained the best mAP (35.28%).
In the multi-camera scenario (without the long-term constraint), performance improves considerably. Here, ResNet152 [17] achieved the highest scores, with 56.01% CMC@1 and 55.90% mAP, demonstrating the advantage of deeper backbones when IDs are observed across multiple views without long interruptions. Although this setting remains challenging due to differences in camera angles and positions, IDs still share more appearance consistency than in the long-term scenarios, which makes this scenario comparatively more feasible.
The reappearance scenario, in contrast, proved the most manageable. ResNet101-IBN achieved the best CMC@1 (65.73%), while ResNet152 obtained the highest mAP (63.85%). These results suggest that reappearances, often within the same camera view, are handled relatively well by current methods compared with the other analyzed scenarios.
Overall, ResNet-based models showed the most consistent and robust performance across all CHIRLA scenarios. While reappearance is comparatively easier, the long-term and multi-camera long-term scenarios remain highly challenging and highlight the need for models that are more resilient to temporal gaps and inter-camera variations.